OpenStudy (anonymous):

find the area of an equalateral triangle, sides are 12. round to the nearest 10th

OpenStudy (anonymous):

First, get height. 6^2 + h^2 = 12^2 so h^2 = 12^2 - 6^2 Then, A = (1/2)bh

OpenStudy (anonymous):

\[Area= \frac{ \sqrt{3}a^2 }{ 4 }\] a= Side

OpenStudy (anonymous):

i have no idea how to do that

OpenStudy (anonymous):

\[\frac{ \sqrt{3}*12*12 }{ 4 } = \sqrt{3}*36\]

OpenStudy (anonymous):

\[\sqrt{3} = `1.73\]

OpenStudy (anonymous):

so 62.28 is the answer ?

OpenStudy (anonymous):

Yup and Round it to Nearest 10 th